OpenStudy (anonymous):

the expression \[b^{-3\2} , b>0 \] is the equivalent to?

OpenStudy (anonymous):

\[=1/b ^{3/2}=1/\sqrt{b^3}\]

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

u welcome